Solve by whatever method seems easiest, quadratic, factoring, property of square, of completing the square.
3x^2 = 2x + 5

Answer by scott8148(6628)

3x^2 - 2x - 5 = 0

factoring ___ (3x - 5)(x + 1) = 0

x = 3/5 ___ and ___ x = -1